OpenStudy (anonymous):

How do you prove the derivative of f(x) = a^x? Attempt: I got to this step: (a^x) * (lim h - > 0 of (a^h - 1/h)) I realized that if I were to sub in f'(0), I would get a^h - h, so, = a^x * f'(0) Now, how do I get from the above to: f'(x) = a^x*ln(a) ?

OpenStudy (anonymous):

proof using limits..

OpenStudy (anonymous):

Yea thats what I started with and then I got this point = a^(x) * f'(0).

OpenStudy (anonymous):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

Ok I am stuck.

OpenStudy (anonymous):

What do I do from where I left off?

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{a^{x+h}-a^{x}}{h}=a^{x}\lim_{h \rightarrow 0}\frac{a^{h}-1}{h}\]So now we have to show that the limit is equal to ln(a).

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Good question. I'm stuck, too.

OpenStudy (anonymous):

The limit as it's written goes to 0/0. I was thinking about using l'Hopital's rule and taking the derivative of top and bottom wrt h, but then I realized that involved using \[\frac{d}{dh}a^{h}\]We would be here forever. I think there must be some substitution we could use, but I don't see what it is.